Find ΔE

   a. when 175 J of work is done on a system that evolves 50 J of heat to the surroundings

   b. when a gas absorbs 28 kJ of heat and has 13 kJ work done on it


1
Expert's answer
2022-03-14T18:17:04-0400

Solution.

"a)\Delta\Epsilon=A-Q;"

"\Delta E=175J-50J=125J;"

"b)\Delta E=Q-A;"

"\Delta E=28kJ-13kJ=15kJ;"

Answer: "a)\Delta E=125J;"

"b)\Delta E=15kJ."